OpenStudy (samsan9):

solve for x. logx+log(x-1)=log(4x)

OpenStudy (aum):

\(\large \log(A) + \log(B) = \log(AB)\) Use the above to combine the two logs on the left hand side. Then use the fact that if log(C) = log(D), then C = D

OpenStudy (samsan9):

log(x) + log(x - 1) = log(4x) log[x(x - 1)] = log(4x) x(x - 1) = 4x x^2 - x - 4x = 0 x^2 - 5x = 0 x(x - 5) = 0 x = 0 or 5 is this right?

OpenStudy (aum):

Correct. But you have to see if there are any extraneous solutions. If we put x = 0 in the original equation we realize it is no good because log is defined only for numbers greater than 0. zero and negative numbers are not allowed for log. So x = 0 is not a valid solution. But x = 5 is a valid solution.

OpenStudy (samsan9):

oh ok so no matter what anything zero or less is wrong ?

OpenStudy (aum):

log(x) is defined only for x > 0.

OpenStudy (aum):

So if you get x = 0 or x = -3 or x = -100 you have to throw it away because log(x) does not exist for those x.